OpenStudy (anonymous):

Let f(x) = 5x - 26 and g(x) = 6x. Part 1 [4 points] What is (f ◦ g)(x) ? Part 2 [2 points] Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

(f ◦ g)(x) = f ( g(x) ) = f( 6x) = 5(6x) - 26 = 30x-26 hope i havent made any mistake

OpenStudy (anonymous):

is that part 1